OpenStudy (anonymous):

HELP PLZZZZZZ...A DVD rental box has a membership plan in which a person pays a flat fee for $10 plus $2.00 for each DVD rented. Non Members pay $4.50 for each DVD rented. How many DVDs would a person have to rent to make a membership plan a better deal?

jimthompson5910 (jim_thompson5910):

let x = number of DVD's rented and y = total cost

jimthompson5910 (jim_thompson5910):

if you have a membership, then it will cost $2 per DVD, so if you rent x of them, then you pay 2x dollars then you add on the $10 flat fee so the total cost equation for members is y = 2x + 10

jimthompson5910 (jim_thompson5910):

For nonmembers, the cost equation is y = 4.50x because it costs $4.50 per DVD for nonmembers

jimthompson5910 (jim_thompson5910):

You have these two equations y = 2x + 10 y = 4.50x find their point of intersection and tell me what you get

OpenStudy (anonymous):

i got the intersection as (4,18) so wht is the meaning of this solution?

OpenStudy (anonymous):

but whts the answer thn?.....more than 4 times?

jimthompson5910 (jim_thompson5910):

That is correct. If you rent more than 4 DVDs, then the membership is the better deal.

OpenStudy (anonymous):

but like, how? to be honest, some1 else told me this answer. and i need to explain/justify my answer. but i dont know how?

jimthompson5910 (jim_thompson5910):

y = 2x + 10 ... start with the first equation 4.50x = 2x + 10 ... plug in y = 4.50x now solve for x

jimthompson5910 (jim_thompson5910):

solving for x gives: 4.50x = 2x + 10 4.50x-2x = 10 2.50x = 10 x = 10/2.50 x = 4

jimthompson5910 (jim_thompson5910):

And then you plug that back into either equation that has x,y in it y = 2x + 10 y = 2*4 + 10 y = 18

jimthompson5910 (jim_thompson5910):

But really all you need is that x value to answer this question

OpenStudy (anonymous):

yes, i know that. but how does this help in finding out on How many DVDs would a person have to rent to make a membership plan a better deal?

jimthompson5910 (jim_thompson5910):

at the point of intersection, the two deals cost the same (for the same number of DVDs rented)

jimthompson5910 (jim_thompson5910):

This point of intersection is (4,18), ie x = 4, y = 18 basically this means that if you rent 4 DVDs, then it will cost $18 whether you are a member or a nonmember

jimthompson5910 (jim_thompson5910):

renting more than 4 DVDs will be cheaper as a member because you're adding on $2 if you are a member (compared to adding on $4.50 for nonmembers)

OpenStudy (anonymous):

Ohh. ok, now I get it. Thank you very much!

jimthompson5910 (jim_thompson5910):

you're welcome